OpenStudy (msmr):

A number x is 2 more than the product of its reciprocal and its additive inverse. In what interval does the number lie?

OpenStudy (anonymous):

\[x+2=\frac{1}{x}\times (-x)\]

OpenStudy (anonymous):

\[x+2=1\] \[x=-1\]

OpenStudy (xishem):

\[x=(-x)(\frac{1}{x})+2\]You have it backwards, I believe, satellite.

OpenStudy (msmr):

would the interval be \[-2<x \le0\]?

OpenStudy (xishem):

\[x=-1+2=1\]1 is two more than -1*(1/1).

OpenStudy (msmr):

oh, if satellite had it backwards that would make sense. The interval would have to be \[0<x \le2\]

OpenStudy (msmr):

right?

OpenStudy (xishem):

The number lies in an infinite number of different intervals. \[-1290 \le x \le 2\]\[0.99999 \le x \le 4\]The question isn't specific to what interval it wants. Your answer works, though.

OpenStudy (msmr):

it's a multiple choice question so it was the only one that worked. thank you! I'm studying for the AMC 10 and it's not fun. :(